OpenStudy (anonymous):

Find the equation of the line that contains the given point and has the given slope. Let y be the dependent variable and let x be the independent variable. Point (4, 5), m = 2/3

OpenStudy (anonymous):

i'll not tell u the answer take help from these examples it will help u more This is a 2 step process. First find the slope of the line thru the points. slope, m = diffy/diffx m = (5-1)/(3-2) m = 4 --------- Now use y = mx + b with either point to find b, the y-intercept. y = mx + b 5 = 4*3 + b b = -7 ------- y = mx + b y = 4x - 7 is the answer. ------------------------ Another approach using determinants: |x y 1| |2 1 1| = 0 |3 5 1| -------- x(1-5) - y(2-3) + 1(10-3) = 0 -4x + y + 7 = 0 y = 4x - 7